OpenStudy (anonymous):

URGENT PLEASE RESPOND. How do I multiply, then simplify -5/6y-4 • -6/5y+6 ?

OpenStudy (shamil98):

\[\frac{ -5 }{ 6y-4 } * \frac{ 6 }{ 5y+6 } = \frac{ -5(6) }{ (6y-4)(5y+4) }\] Simplify the denominator using FOIL, the numerator is just constants so that is pretty easy to simplify.

OpenStudy (anonymous):

Okay thank you very much !! @shamil98